OpenStudy (anonymous):

Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3

OpenStudy (anonymous):

i got 63

OpenStudy (anonymous):

it could also be 15

OpenStudy (gudden):

-3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3 lets try .. if we could simplify the equation further and yes we can... -3a2 - b3 + 3c2 - 2b3 and since we could see b^3 twice terefore lets make it once so now the equation becomes -3a^2 -3b^3 +3c^2 now even take the three common we get 3 ( -a^2 - b^3 + c^2) now using the value of a as a = 2 we get 3 ( -(2)^2 - b^3 + c^2) i.e. 3( -4 - b^3 + c^2) and now using b = -1 we get 3( -4 - (-1) + c^2) now using c = 3 we get 3( -4 +1 + 9) which is equals to 3(6) = 18...!